What is the pH of a solution that has a hydroxide concentration of 1.9 X 10-5...

  1. What is the pH of a solution that has a hydroxide concentration of 1.9 X 10-5 M?

  2. What is the pH of a 0.056 M solution of H2SO4? H2SO4 is a strong acid.

  3. Calculate the pH of a 0.50 M solution of Sr(OH)2. Sr(OH)2 is a strong base.

  4. Mylanta liquid has 200 mg of Al(OH)3 and 200 mg of Mg(OH)2 per teaspoonful. Write the equation for the neutralization of stomach acid. Assuming both bases are completely soluble in water, calculate the mass of stomach acid that can be neutralized with one teaspoon of Mylanta.

  5. Lye has 400 mg of NaOH per teaspoon. Calculate the mass of battery acid that can be neutralized by 1 teaspoon of lye. Assume that battery acid is only 5% H2SO4 by mass.
Answer #1

Given,

[OH-] = 1.9 x 10-5 M

Autoionization constant of water, Kw = 10-14 at 25 oC

Now,

Kw = [H3O+][OH-] = 10-14

or, [H3O+] = 10-14/(1.9 x 10-5)

                 = 5.3 x 10-10 M

Hence, the concentration of H3O+ = 5.3 x 10-10 M

Now,

pH = -log[H3O+]

      = -log(5.3 x 10-10)

      = 9.3

Hence, the pH of the solution = 9.3